Convert:

1) 3.4 x 10⁴ BTU to Kj

2) 3472 BTU/IBM to kj/kg

3) 70ibm/ft to kg/me

4) 90°F to k

5) 120btu/hr ft² to j/sm²


1
Expert's answer
2021-02-01T15:07:39-0500

"1) \ \ 3.4\cdot 10^4\ BTU=35872\ kJ"


"2)\ \ 3472\ BTU/lbm=8076\ kJ/kg"


"3)\ \ 70\ lbm/ft=104.2\ kg/m"


"4)\ \ 90°F=32.2°С"


"5)\ \ 120\ BTU/hr ft^2=1362784.6\ J/(s\cdot m^2)"